a+2b+3c=0 eq -3a+2b+5c=0 find the value of a b and c

Grade:12

1 Answers

Vikas TU
6 years ago
Dear Student,
given equations: a+2b+3c=0 and-3a+2b+5c=0
so solving them,
(a+2b+3c=0)*3
+ (-3a+2b+5c)=0 
we h=get b=-7c/4 and c=c
putting value of b and c in a+2b+3c=0,
we get a+2*(-7c/4) +3c=0
=>a=c/2
so c=c , a=c/2 and b=-7c/4 [Ans]
 
Cheers!!
Regards,
Vikas (B. Tech. 4th year
Thapar University)